# DexCatch: Learning to Catch Arbitrary Objects with Dexterous Hands

### v1

@@ Abstract
Achieving human-like dexterous manipulation remains a crucial area of research in robotics. Current research focuses on improving the success rate of pick-and-place tasks. Compared with pick-and-place, throw-catching behavior has the potential to increase picking speed without transporting objects to their destination. However, dynamic dexterous manipulation poses a major challenge for stable control due to a large number of dynamic contacts. In this paper, we propose a Stability-Constrained Reinforcement Learning (SCRL) algorithm to learn to catch diverse objects with dexterous hands. The SCRL algorithm outperforms baselines by a large margin, and the learned policies show strong zero-shot transfer performance on unseen objects. Remarkably, even though the object in a hand facing sideward is extremely unstable due to the lack of support from the palm, our method can still achieve a high level of success in the most challenging task. Video demonstrations of learned behaviors and the code can be found on the supplementary website.
I INTRODUCTION
The study of dexterous hands is a widely pursued research area aimed at matching the intricate capabilities of human hands. Researchers aspire to employ these hands for delicate manipulations [1, 2, 3, 4]. In contrast to traditional planning approaches [5], reinforcement learning has witnessed remarkable success in robot planning tasks [6, 7, 8, 9]. Consequently, researchers have applied reinforcement learning techniques to various tasks involving dexterous hands, leading to diverse outcomes [10, 11, 12]. Previous work has tackled a series of manipulation tasks, including in-hand reorientation [11, 13], opening a door [10], assembling LEGO blocks [14], and even solving a Rubik’s cube [12]. These tasks can be summarized as common behaviors in daily living, where hands pick up an object, reorient it in hand, and then either place it or use it as a tool [13].
While pick-and-place can fulfill the requirements of many daily tasks, dynamic throw-catching holds the potential for more efficient task execution [15, 16, 17]. For example, two robots can employ throwing and catching to efficiently transfer objects instead of pick-and-place, as shown in Fig. 1. Meanwhile, it can extend the working space of robots, which avoids the potential collisions for robot interactions [15]. However, precisely catching arbitrary objects proves to be highly challenging [18]. First, the dynamic contact with strong force makes the object fall from the hand. In previous work [18], model-free reinforcement learning methods achieved nearly 0% success rate in the throwing-catching task within 30 million environment steps. Secondly, varying object-centric properties (e.g. mass distribution, friction, shape) require the catching behavior robust enough, especially for objects of daily life. Thirdly, faced with flying objects from random directions, the catching hand should use different postures. But the challenge is the task of catching objects with hands facing sideward, where objects easily fall from the hands due to the lack of any support surface, as illustrated in Fig. 1.
To address the challenge of dexterous two-handed throwing and catching, we present the Stability-Constrained Reinforcement Learning (SCRL) algorithm. This algorithm introduces three enhancements that significantly boost the exploration of stable catching behavior. 1) Stability Condition: The algorithm enforces stability in the state during exploration, enhancing the strategy’s robustness. 2) Intrinsic Advantage: Strategies are encouraged to explore actions with high advantage, contributing to more effective exploration. 3) Normalization Method: By constraining parameter representation space, the learning process is accelerated. Our method outperforms baselines on various throwing and catching scenarios where the hands face upward and sideward. Furthermore, the learned policy generalizes to create diverse catching gestures on the objects with varying properties (e.g., mass distribution, friction, shape). It is worth noting that it achieves a success rate for unseen objects, even on the most challenging catching task with the hands facing sideward.
@@ I INTRODUCTION
Figure 1: Generalization tasks for throwing and catching objects with shadow hands. Shadow hands can throw and catch a variety of objects with different shapes, masses and other properties when the hands face sideward.
II Related Work
@@
The versatile dexterous hand plays a crucial role in performing various tasks across different human-focused settings. However, effectively managing dexterous hand systems poses a challenge due to their complex actions and intricate contact interactions [5]. Traditionally, controllers for manipulation tasks heavily rely on dynamic models and trajectory optimization techniques [1, 2, 3, 4]. This involves employing simplified dynamic models to generate control patterns using trajectory planning approaches. For instance, Williams et al. [19] achieved success using the Model Prediction Path Integration Control (MPPI) method to skillfully manipulate a cube. Charles Voss et al. [18] enhanced the MPPI method, making it simpler to manage task handovers between two hands.
@@
In recent times, there has been a notable surge in the adoption of controller designs rooted in reinforcement learning. These approaches simplify the controller design process, and model-agnostic techniques have gained substantial popularity within the control domain. In the realm of dexterous manipulation, numerous endeavours have showcased remarkable advancements compared to conventional controllers [20, 21, 22]. Entities like OpenAI have harnessed reinforcement learning-based controllers to successfully rearrange blocks and solve Rubik’s cubes [12]. To expedite training, a blend of reinforcement learning and imitation learning has been employed to facilitate dexterous hands in acquiring diverse skills such as repositioning objects, utilizing tools, opening doors, and more [10]. Acknowledging the existing method’s limited ability to generalize, Chen et al. [13] introduced an effective system capable of learning how to reorient a wide array of objects. For the cooperation between two hands, Zakka et al. [23] and Chen et al. [22] proposed a series of challenging tasks for bi-manual dexterity, like mastering the piano and lifting the pot. Furthermore, a recent work introduced the Population-Based Training (PBT) algorithm in RL learning, leading to solving the two-handed reorientation task [21]. However, the work only focuses on a single object. Our method successfully achieves the throwing and catching of diverse objects, showcasing superior performance compared to baseline approaches.
III Dexterous Catching Tasks
III-A Preliminaries
@@
The throwing and catching tasks of dexterous hands can be formalized as an infinite horizon Markov Decision Process (MDP), which is defined by a tuple $(S,A,R,P,\rho_{0})$ with the transition function $P(.|s,a)$. Among them, $S\in R^{n}$ is the current state space, $A\in R^{m}$ is the current action space, $R$ is the reward, $\gamma$ $\subset[0,1)$ is the discounted factor, and $\rho$ is the initial state distribution. $S^{\prime}$ and $A^{\prime}$ is the next state space and the next action space. The goal of RL is to find an optimal cumulative return R of the MDP with the policy $\pi:S\leftarrow A$. The problem is shown to be represented as an optimization problem of the expected discounted cumulative return.
$\mathbb{E}_{s_{0}\sim\rho_{0},a_{t}\sim\pi(\cdotp|s_{t}),s_{t+1}\sim p(\cdotp|s_{t},a_{t})}[\sum_{i=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$ (1)
Figure 2: The features of different point clouds are extracted by PointClipV2 [24], then reduced by PCA, and finally visualized in two dimensions. The Clustering Using Representatives(CURE) method shows that any object is approximately divided into three categories according to different shapes, extending in three directions, respectively, elongated, flat, and regular. Sample of objects each category are bottles, keyboards, and cups, respectively.
Figure 3: A: The SCRL algorithm takes as input the environmental observation and the point cloud feature of the object. Then, it learns the catching policy for dexterous hands through an Actor-Critic structure. B: The Lyapunov function, the policy function and the value function are estimated using neural networks. Their network structures are similar, and the only difference is the dimension of the output. For example, the Lyapunov function contains a linear layer, a batch normalization layer, and a penultimate normalization layer, and the last linear layer takes a scalar as output.
III-B Task Desciption
@@
Our research centers on the challenges of throwing and catching tasks using dexterous shadow hands. Our aim is for these shadow hands to adeptly handle diverse objects in various positions and execute agile catches using different gestures. To address this objective, our design comprises of three key components.
@@
The shadow hand models, which emulate human skeletons, possess 24 degrees of freedom (DOFs) and are operated using 20 pairs of agonist-antagonist tendons [13]. Meanwhile, the remaining four joints are underactuated. The presence of force sensors on the ten fingertips facilitates the acquisition of force-related data during catching and throwing maneuvers. In our paper, we investigate three catching scenarios, namely Abreast Catch, Underarm Catch and Overarm Catch, which are shown in Fig. 5 respectively.
In Abreast/Underarm Catch, a hand facing upward should throw an object to the other hand, and the other one should hold it in the palm later. Notably, Overarm Catch also solves the same problem but represents the most challenging catching task. This is because the object in the hands is not stable without the support from the palm when the hands are oriented vertically.
@@
Our design objective mainly involves attaining the capability to generalize across diverse objects through the introduction of point cloud features of objects. To achieve this, we leverage the IsaacGym simulator [25], which enables parallel training, allowing the system to learn the skill of catching various objects simultaneously.

### v2

@@ Abstract
Achieving human-like dexterous manipulation remains a crucial area of research in robotics. Current research focuses on improving the success rate of pick-and-place tasks. Compared with pick-and-place, throwing-catching behavior has the potential to increase the speed of transporting objects to their destination. However, dynamic dexterous manipulation poses a major challenge for stable control due to a large number of dynamic contacts. In this paper, we propose a Learning-based framework for Throwing-Catching tasks using dexterous hands (LTC). Our method, LTC, achieves a 73% success rate across 45 scenarios (diverse hand poses and objects), and the learned policies demonstrate strong zero-shot transfer performance on unseen objects. Additionally, in tasks where the object in hand faces sideways, an extremely unstable scenario due to the lack of support from the palm, all baselines fail, while our method still achieves a success rate of over 60%. Video demonstrations of learned behaviors and the code can be found on the supplementary website.
Keywords: Reinforcement Learning, Dexterous Manipulation, System Stability
@@
Figure 1: Generalization tasks for throwing and catching objects with shadow hands. Shadow hands can throw and catch a variety of objects with different shapes, masses, and other properties when the hands face sideward.
1 Introduction
The study of dexterous hands is a widely pursued research area aimed at matching the intricate capabilities of human hands. Researchers employ these hands for delicate manipulations [1, 2, 3, 4]. Recently, researchers have applied reinforcement learning (RL) techniques to various tasks involving dexterous hands, leading to diverse outcomes [5, 6, 7]. They have tackled a series of manipulation tasks, including in-hand reorientation [6, 8], opening a door [5], assembling LEGO blocks [9], and even solving a Rubik’s cube [7]. These tasks can be summarized as common behaviors in daily living, where hands pick up an object, reorient it in hand, and then either place it or use it as a tool [8].
While pick-and-place can fulfill the requirements of many daily tasks, dynamic throwing-catching holds the potential for more efficient task execution [10, 11, 12]. For example, two robots can employ throwing and catching to efficiently transfer objects instead of pick-and-place, as shown in Fig.1. Meanwhile, it can extend the working space of robots, which avoids the potential collisions for robot interactions [10]. However, precisely catching arbitrary objects proves to be highly challenging [13]. 1) First, the dynamic contact with strong force makes the object fall from the hand. In previous work [13], model-free reinforcement learning methods achieved nearly 0% success rate in the throwing-catching task within 30 million environment steps. 2) Secondly, varying object-centric properties (e.g. mass distribution, friction, shape) require the catching behavior robust enough, especially for objects of daily life. To target this problem, previous methods must learn the dynamics of a moving object [14]. The additional module requires expert demonstrations and increases the system’s complexity. 3) Thirdly, when faced with flying objects coming from random directions, the catching hand must adopt different postures, enabling a larger working space. However, the challenge arises in catching objects with hands facing sideways, where objects easily fall due to the lack of any supporting surface, as illustrated in Fig. 1. Due to this challenge, the task of catching objects with hands facing sideways has not been explored previously.
To address the aforementioned challenges, we propose a Learning-based framework for Throwing-Catching tasks using dexterous hands (LTC). The simple but effective framework can achieve superior performance across diverse daily objects, in various throwing and catching scenarios where the hands face upward or sideward. The main contributions of this work are as follows.
  • We propose a simple and effective framework using a model-free reinforcement learning algorithm, to solve the throwing-catching tasks. Particularly, it can catch the flying objects even on the most challenging catching task with the hands facing sideward.
  • We propose a hybrid advantage estimation method, which considers Lyapunov stability. While the effectiveness of the stability condition has primarily been verified for robustness, this is the first time it has been observed to benefit the generation of stable catching behavior for flying objects, thereby improving the accuracy of throwing-catching.
  • By using compressed point cloud features of objects as input and applying domain randomization during training, our method excels in generalizing to daily objects with diverse properties (e.g., mass distribution, friction, shape, and initial pose), and performs effectively even on previously unseen objects.
2 Related Work
@@
The versatile dexterous hand plays a crucial role in performing various tasks across different human-focused settings. However, effectively managing dexterous hand systems poses a challenge due to their complex actions and intricate contact interactions [15]. Traditionally, controllers for manipulation tasks heavily rely on dynamic models and trajectory optimization techniques [1, 2, 3, 4]. The paper “Catching Objects in Flight” is highly regarded in this field [14]. To enhance model accuracy, their method learns the dynamics of a moving object and develops a reachable and graspable model using concise demonstrations. Furthermore, Williams et al. [16] achieved success using the Model Prediction Path Integration Control (MPPI) method to skillfully manipulate a cube. Charles Voss et al. [13] improved the MPPI method to manage task handovers between two hands. However, they focused only on a single object due to the complexities involved in model learning.
@@
In recent times, RL-based approaches simplify the controller design process, and model-agnostic techniques have gained substantial popularity within the control domain. In the realm of dexterous manipulation, numerous endeavors have showcased remarkable advancements compared to conventional controllers [17, 18, 19]. Entities like OpenAI have harnessed reinforcement learning-based controllers to successfully rearrange blocks and solve Rubik’s cubes [7]. To expedite training, a blend of reinforcement learning and imitation learning has been employed to facilitate dexterous hands in acquiring diverse skills such as repositioning objects, utilizing tools, opening doors, and more [5]. Acknowledging the existing method’s limited ability to generalize, Chen et al. [8] introduced an effective system capable of learning how to reorient a wide array of objects. For the cooperation between two hands, Zakka et al. [20] and Chen et al. [19] proposed a series of challenging tasks for bi-manual dexterity, like mastering the piano and lifting the pot. Furthermore, a recent work introduced the Population-Based Training (PBT) algorithm in RL learning, leading to solving the two-handed reorientation task [18]. However, most recent works [18, 21] only focus on a single object or a kind of posture of hands. Our method successfully achieves the throwing and catching of diverse objects, especially for the task with hands facing sideward.
3 Preliminaries
3.1 Reinforcement Learning
@@
The throwing-catching tasks of dexterous hands can be formalized as an infinite horizon Markov Decision Process (MDP), which is defined by a tuple $(S,A,p,R,\rho_{0},\gamma)$. Among them, $S\in\mathbb{R}^{n}$ is the state space, $A\in\mathbb{R}^{m}$ is the action space, $p:S\times A\times S\rightarrow[0,1]$ is the transition probability distribution, $r:S\times A\rightarrow\mathbb{R}$ is the reward, $\gamma$ $\in[0,1)$ is the discount factor, and $\rho_{0}:S\rightarrow[0,1]$ is the initial state distribution. The goal of RL is to find an optimal cumulative return of the MDP with the policy $\pi:S\times A\rightarrow[0,1]$. The problem is formulated as maximizing the expected discounted cumulative return: $\mathbb{E}_{s_{0}\sim\rho_{0},a_{t}\sim\pi(\cdotp|s_{t}),s_{t+1}\sim p(\cdotp|s_{t},a_{t})}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$.
3.2 Proximal Policy Optimization(PPO)
[…]
$L(\cdotp)=\min\left\{\begin{aligned} &\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})}A^{\pi_{\theta_{k}}}(s_{t},a_{t}),&\mathrm{clip}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right]A^{\pi_{\theta_{k}}}(s_{t},a_{t})\end{aligned}\right\}$ (1)
where $\epsilon$ is a hyperparameter, ${A^{\pi_{\theta_{k}}}(s_{t},a_{t})}=G_{t}^{\pi_{\theta_{k}}}-V^{\pi_{\theta_{k}}}_{\phi}(s_{t})$, is the advantage function. $G_{t}^{\pi_{\theta_{k}}}$ is cumulative returns. For parameters $\phi$, the update is expressed as:
$\begin{array}[]{l}\phi_{k+1}=\underset{\phi}{\arg\min}\underset{s_{t},a_{t}\sim\pi_{\theta_{k}}}{\mathbb{E}}[(G_{t}^{\pi_{\theta_{k}}}-V^{\pi_{\theta_{k}}}_{\phi}(s_{t}))^{2}]\\ \end{array}$ (2)
However, because the calculation of $G_{t}^{\pi_{\theta_{k}}}$ value needs the whole data of a finished episode, we approximate $G_{t}^{\pi_{\theta_{k}}}$ to $r_{t}+\gamma V^{\pi_{\theta_{k}}}_{\phi}(s_{t+1})$ through TD learning method.
4 Dexterous Catching Tasks
@@
Our research centers on the challenges of throwing-catching tasks using dexterous shadow hands. Our aim is for these shadow hands to adeptly handle diverse objects in various positions and execute agile catches using different gestures. To address this objective, our design comprises three key components.
@@
The shadow hand models, which emulate human skeletons, possess 24 degrees of freedom (DOFs) and are operated using 20 pairs of agonist-antagonist tendons [8]. The presence of force sensors on the ten fingertips facilitates the acquisition of force-related data during catching and throwing maneuvers. In our paper, we investigate five catching scenarios, namely Overarm Catch, Overarm2Abreast Catch, Under2Overarm Catch, Abreast Catch, Underarm Catch and which are shown in Fig. 4 respectively. In these five tasks, one hand should throw an object to the other hand, and the other one should hold it in the palm later. Some tasks with vertical hands are challenging. It is because the object in the hands is not stable without the support from the palm.
@@
Our design objective is to generalize across diverse objects by introducing point cloud features, which are obtained from a pre-trained model, PointClipV2 [22]. Additionally, we apply Principal Component Analysis (PCA) to compress the features. The experiments illustrate that the shapes of different objects can be distinguished using the compressed features. More details can be found in Appendix B.
State, Action and Reward
1) State: The state space consists of 424 elements. Among these, the 422-dimensional state incorporates crucial information: hand position, hand orientation, angular and linear velocities of the hands, position and speed of each hand’s degree of freedom, force-torque sensor readings of the hands, target object’s position and orientation, as well as the current object’s position and orientation. The above observation corresponds to the setting in [19]. Additionally, a two-dimensional state features the input derived from the point cloud data of the object. It serves as the perceptual part for the agent regarding the object. 2) Action: The action space comprises 52 dimensions, encompassing 20 pairs of agonist-antagonist tendons for both hands, along with three dimensions each for the required for ce and torque of each hand. 3) Reward: The reward design comprises three components: the difference in position and posture between the object and the target, the penalty for the action scale of the shadow hand joints, and a combination of task completion bonuses and failure costs. We explain the details of the reward in Appendix C.
Figure 2: A: The method takes as input the environmental observation and the point cloud feature of the object. Then, it learns the catching policy for dexterous hands through an Actor-Critic structure. B: The Lyapunov function, the policy function and the value function are estimated using neural networks. Their network structures are similar, and the only difference is the dimension of the output. For example, the Lyapunov function contains a linear layer, a batch normalization layer, and a penultimate normalization layer, and the last linear layer takes a scalar as output.
5 Method
In this section, we introduce the LTC algorithm in detail. LTC is an end-to-end framework based on the Proximal Policy Optimization (PPO) [23] algorithm. Since the throwing-catching task is a dynamic and agile behaviour, the basic PPO algorithm encounters two key challenges: How to achieve efficient throwing and stable catching without falling? We make some algorithmic modifications to the basic PPO algorithm. First, to accelerate the learning of throwing, we design an intrinsic advantage and add penultimate normalization layers in the network. Secondly, a stable catching behaviour is holding the object in the palm without falling the object. As depicted in Fig. 3, the maximum sum reward (optimality) can not guarantee the generation of the stable catching behaviour. To encourage more stable catching, we include the Lyapunov stability condition in policy learning. The specific formulation can be found in Section 5.2.1.
5.1 Framework
Figure 3: To enhance the success rate, our approach prioritizes optimizing rewards while ensuring the stability of the grasping gesture. By integrating Lyapunov stability alongside reward optimization, we mitigate the risk of completing tasks in unstable poses. The size of the stable region varies based on task difficulty, with simpler tasks offering a larger stable region.
The architecture is illustrated in Fig. 2. Here’s a breakdown of its components:1) Point Cloud Feature Extraction: The initial step involves extracting point cloud features using the PointClipV2 [22]. Subsequently, Principal Component Analysis (PCA) is employed to reduce dimensionality, yielding a 2-dimensional feature output. This output serves as the perception state for the shadow hands system; 2) Observation Input: The agent takes both the current observation of objects and the shadow hands and the compressed features from point clouds as input; 3) Actor-Critic Framework: The actor and critic are represented as two neural networks, which is the same as PPO algorithm. The network architecture for the Lyapunov function is depicted on the right side of Fig. 2. This structure comprises three linear layers, with a Batch normalization layer and an ELU activation function appended to the output of each layer. Following these, there is a penultimate normalization layer [24], succeeded by a linear layer to regulate the output dimension. The network structure for the policy and value function are similar to that of the Lyapunov value network.
5.2 Hybrid Advantage Estimation
The hybrid advantage $A_{all}$ consists of 3 parts: standard advantage estimation based on PPO, stability advantage estimation, and intrinsic advantage estimation. The formula for the hybrid advantage is expressed as follows:
$\begin{array}[]{cc}A_{all}(s_{t},a_{t},s_{t+1})=\beta_{1}A^{\pi_{\theta}}(s_{t},a_{t})+\beta_{2}A_{L}(s_{t},s_{t+1})+\beta_{3}A_{I}(s_{t},a_{t})\\ \end{array}$ (3)
$\beta_{1}+\beta_{2}+\beta_{3}=1$ (4)
where $A^{\pi_{\theta}}(s_{t},a_{t})$ refers to the time differential advantage function of the PPO algorithm, $A_{L}(s_{t},a_{t})$ represents the Lyapunov stability advantage function, $A_{I}(s_{t},s_{t+1})$ represents the intrinsic advantage function, and $\beta_{1},\beta_{2},\beta_{3}\in[0,1]$ are hyperparameters. As follows, we introduce the Lyapunov stability advantage and intrinsic advantage in detail.
5.2.1 Lyapunov Stability
Figure 4: Video snapshots of throwing-catching tasks performed by our method. There exist five tasks with different postures of hands, in order from difficult to easy: Overarm Catch, Overarm2Abreast Catch, Under2Overarm Catch, Abreast Catch, and Underarm Catch. The most challenging task is Overarm Catch because the hands are both oriented vertically. Notably, our method can catch diverse objects in five tasks.
The Lyapunov function serves as a potent tool for assessing system stability. One approach, known as the self-learning Lyapunov stability method [25], is utilized in our paper. First, we introduce a Lyapunov stability critic network, which represents the candidate Lyapunov function $V_{L}$ using neural networks. Subsequently, the empirical Lyapunov risk $\mathcal{R}(\Lambda)$ [25] is defined as per Eq. (5). The three parts in Eq. (5) represent three requirements for the Lyapunov function [25]. The parameter update is then executed by minimizing the loss function associated with the Lyapunov risk $\mathcal{R}(\Lambda)$, as part of achieving the desired system stability.
$\mathcal{R}(\Lambda)=\mathbb{E}[\max(-V_{L}(s_{t}),0)+\max(0,L_{f}V_{L}(s_{t}))+V_{L}^{2}(0)]$ (5)
where $f$ is the dynamical system, $s_{t}$ is the state at step $t$. The empirical Lyapunov risk is non-negative, we approximate the Lie derivative $L_{f}V_{L}(s_{t})$ by finite difference of the system’s sampling trajectory because of the complexity of modeling this system’s dynamics:
$\begin{array}[]{cc}L_{f}V_{L}(s_{t})=\frac{1}{\Delta t}(V_{L}(s_{t+1})-V_{L}(s_{t}))\end{array}$ (6)
where $\Delta t$ is the time difference between them. Since the Lyapunov function is considered to be another value function, the Lie derivative term $L_{f}V_{L}(s_{t})$ is the only term in the Lyapunov risk affected by the policy. Therefore, an additional goal of ensuring $L_{f}V_{L}(s_{t})<0$ is added to the policy optimization, to encourage the policy’s update with the requirement of stabilization. However, the classical constraint $L_{f}V_{L}(s_{t})<0$ makes the value of the Lyapunov function likely to be stuck in a minimal number close to 0. This drawback hinders the process of policy improvement due to the lack of an explicit gradient.
Therefore, we design a more strict constraint $L_{f_{\pi_{\theta}}}V_{L}(s_{t})<-kV^{\alpha}_{L}(s_{t})$, where $\alpha$ and $k$ are constants ranging from 0 to 1. Note that $kV^{\alpha}_{L}(s_{t})$ is non-negative. When $kV^{\alpha}_{L}(s_{t})$ is larger, an intuitive phenomenon is that $V_{L}$ approaches 0 faster. In other words, the system can reach a stable region in fewer steps, which benefits the generation of stable catching behaviors. More details can be found in Appendix D. Hereafter, we design the part of the Lyapunov stability advantage function based on the clipped Lie derivative term:
$A_{L}(s_{t},a_{t})=\min(kV^{\alpha}_{L}(s_{t}),-L_{f_{\pi_{\theta}}}V_{L}(s_{t}))$ (7)
5.2.2 Intrinsic Advantage
Based on the advantage function of the PPO algorithm, in order to improve the ability of the system efficiently, we design the intrinsic advantage function with reference to the intrinsic reward mechanism [26]. For a Markov decision process trajectory, different values $V_{\phi}(s_{t})$ will affect the success of the final result of this trajectory. To provide a specific example, let’s consider the fact that the next states $s_{t+1}$ exhibit varying conditions, leading to different values $V_{\phi}(s_{t+1})$. It’s worth noting that in a successful process, the value $V_{\phi}(s_{t+1})$ tends to be higher. Consequently, within a Markov decision process, we can further refine the advantage function $A(s,a)$ using empirical measures. If an action is effective, it should result in a higher value, as represented by $V_{\phi}(s_{t+1})-V_{\phi}(s_{t})>0$. This signifies that action $a_{t}$ transitions to a state with a higher success rate. Conversely, an unfavorable action can lead to a decrease in value, reflected as $V_{\phi}(s_{t+1})-V_{\phi}(s_{t})<0$. In order to reduce failure in the early stage of the throwing process, we consider the intrinsic advantage $A_{I}(s_{t},s_{t+1})$ as part of the total advantage function to accelerate the RL training, shown as follows:
where $k$ is a positive constant.
6 Experiments
To test the algorithm’s ability to generalize, we train and test it with objects from the YCB__\__Pybullet dataset [27]. The results show that our algorithm outperforms baseline methods. Four essential components of our algorithm improve its performance. Additionally, our method showcases the ability to generalize different gestures across five skilful throwing-catching tasks, effectively allowing for the successful manipulation of various objects. Importantly, our approach extends its capabilities to catch unseen objects successfully.
Figure 5: The average reward curve for training multiple objects for five different tasks. The first row is the comparison experiment reward graph. The second row is an ablation experiment reward graph. (Curves smoothed report the mean 5 times, and shadow areas show the variance.)
6.1 Comparison Experiments
In this section, we conduct a comparative analysis of our algorithm against mainstream reinforcement learning (RL) algorithms in dexterous hands’ throwing-catching tasks. These algorithms include TRPO [28], SAC [29], DDPG [30], and PPO [23]. The results, as illustrated in Fig. 5, are quite telling. In the most challenging task, the Overarm Catch, our algorithm significantly outperforms the baseline methods. For the Over2abreast Catch, Under2Overarm Catch, and Abreast Catch tasks, our approach demonstrates superior learning efficiency in acquiring throwing and catching skills. Furthermore, even in the relatively easier task, Underarm Catch, our method exhibits slightly improved learning efficiency and performance compared to the baseline algorithms. However, the variance of the training curve is larger than those baselines in Underarm Catch. This is because simpler tasks have more gestures to complete tasks (see Fig. 3), leading to numerous feasible solutions. Thus, the strong exploration of our method results in a high variance of the sum of rewards. Conversely, complex tasks feature limited gestures, resulting in minimal reward variance post-completion.
Video snapshots illustrating the performance in five tasks can be seen in Fig. 4. It’s worth emphasizing that we train the model using nine objects with diverse shapes and mass distributions. This showcases the algorithm’s robust generalization ability, as it can adapt to various objects and perform tasks efficiently and reliably.
6.2 Ablation Experiments
We evaluate the impact of four crucial components in our algorithm. To clarify, here are their definitions: a. indicates the addition of a Lyapunov stability component. b. indicates the addition of an intrinsic advantage component. c. indicates network structure parameter normalization, and d. indicates that the compressed point cloud feature of the object is added as a prior to the state. The effect of these components is shown in the reward curve in Fig. 5. It can be seen from the figure that each component of our method benefits the performance compared with the PPO algorithm. The inclusion of Lyapunov stability (a.) in the baseline PPO notably enhances the hand’s catching ability, particularly accentuated during complex tasks like the Overarm Catch. There exists a more detailed explanation of ablation experiments in Appendix E.
6.3 Generalization Ability
Table 1: Success rate of Task
[…]
Train (9 objects) Test (11 objects)
META Apple(regular) Banana(elongated) Poker(flat) META Stapler Scissor Washer Bowl
Overarm 60.30 63.62 73.51 73.53 67.58 69.71 56.83 76.00 77.34
Overarm2Abreast 68.77 79.01 82.89 68.76 76.79 81.24 83.90 67.42 86.96
Under2Overarm 73.95 75.35 82.21 78.79 73.41 76.85 68.46 57.34 81.88
Abreast 72.95 67.41 88.65 86.34 77.89 94.84 57.52 83.94 83.59
Underarm 88.99 95.87 86.50 90.57 80.47 84.10 63.63 63.10 83.23
  • 1
    After dividing the data into training and test sets, we perform distinct success rate tests(10k times) for single objects and multiple objects. Specifically, we only display the success rate of three typical objects within the training set and four typical objects within the test set. We utilize a reward threshold to ascertain the success rate. In the Under2Overarm Catch task, success is determined if the reward surpasses 20 because of the longer execution process, while for other tasks, a reward exceeding 15 indicates success.
We evaluate our method in the throwing-catching task using various objects, quantifying the success rate as a key metric. We run the testing task in 10k episodes and use the average reward threshold to determine whether it is successful. Since our algorithm pays more attention to the efficiency and stability of generalization ability learning, we no longer use the distance threshold of the final state from the target position as the success rate. The success rate of different objects is illustrated in the Table 1. Among them, we use 11 unseen objects (test set) to verify the generalization ability of our method. Specifically, we adopt two methods to test objects in the training and test sets: i. Test the success rate of multiple objects in parallel. ii. Test the success rate of each object individually. We take the most challenging task, Overarm Catch as an example. The average success rate for the training set objects reaches $60.30\%$, and the highest of individual tests on each object was pen up to $81.23\%$. In addition, the average success rate of the test set object is close to $67.58\%$. Most importantly, our method obtains a good success rate, $56.83\%$, in testing scissors, which is a new but difficult object. Furthermore, we can observe that our method can provide a robust catching strategy because there is not a significant gap between the success rates of training and test objects. The results provide strong evidence of our algorithm’s robust generalization capability for catching diverse objects. We perform a more detailed analysis of the generalization ability in Appendix E.
7 Conclusion
Summary: In this study, we introduce an end-to-end framework (LTC) based on model-free reinforcement learning to tackle throwing-catching tasks with dexterous hands. Remarkably, our method demonstrates exceptional generalization ability, even for unseen objects, and effectively handles scenarios where the hands face sideways or other postures. This research underscores the potential of RL-based approaches in improving object transportation efficiency through dynamic manipulation with dexterous hands.
Limitation: Addressing the randomness for more complex objects or the initial position and posture of the hands is a key challenge for our method. To enhance robustness and success rate, we plan to iteratively refine task randomness by combining curriculum learning methods. Furthermore, we intend to optimize the motion and dynamic control parameters through adaptive parameter learning methods. Additionally, we’ve constructed a robot hardware platform comprising a mobile chassis, a 7-DoF robotic arm, and an allegro hand. Through the meticulous design of domain randomization terms, we hope to achieve zero-shot simulation-to-real transfer and execute throwing-catching tasks in real-world settings.
Appendix A Implementation detail
In our experiments, we conduct training and testing on a platform equipped with an RTX A6000 GPU and an AMD CPU. We utilize nine objects for training across five different task types and test the algorithm with four objects. Our approach also leverages meta-learning for multi-object parallel training. The objects used in our experiments are sourced from the YCB dataset [27]. Due to the limitations of the platform, we employ 64 parallel environments per object for the Abreast Catch task and 256 parallel environments for others. The policy updates every 8 environment steps, and the length of each episode is set to 100 steps. Furthermore, every experiment is built using early reset functions mentioned in Appendix B. To determine the reward threshold, we randomly sampled 50 experimental environments during the testing phase. By visualizing and analyzing the corresponding reward values after successfully catching the objects, we calculated the average and rounded it to establish the reward threshold.
Appendix B Object-generalized design
Figure 6: The features of different point clouds are extracted by PointClipV2 [22], then reduced by PCA, and finally visualized in two dimensions. The Clustering Using Representatives(CURE) method shows that any object is approximately divided into three categories according to different shapes, extending in three directions, respectively, elongated, flat, and regular. Sample of objects in each category are bottles, keyboards, and cups, respectively.